OpenStudy (scarlettfarra2000):

I need help doing the last part of this equation could someone please help me? Write the following equation in standard form. Then solve

OpenStudy (scarlettfarra2000):

\[6q^2-2q=5q-6q+21\]

OpenStudy (scarlettfarra2000):

The standard form of this quadratic equation is \[q^2+4q-21=0\]

OpenStudy (scarlettfarra2000):

How would I get \[(q+?)(q-?)\]

OpenStudy (scarlettfarra2000):

@dayakar Do you want me to write out how I got my answer?

OpenStudy (dayakar):

\[q ^{2}+7q-3q-21=0\] spilitting the middle term to get the factors q*q +7q - 3q - 3*7=0 q(q+7) -3(q+7) =0 (q+7) (q-3) =0

OpenStudy (dayakar):

do u understand the solution

OpenStudy (triciaal):